### v1

@@ Appendix G Proof of Theorem 14
$\mu_{\beta}^{\gamma_{s}}(B_{M+1})=\mathbb{P}\left(|X|\geq(CM)^{1/m}\right)\leq\frac{\mathbb{E}[|X|^{m}]}{\left((CM)^{1/m}\right)^{m}}\leq\frac{1}{M},$ (59)
@@
$\int_{B_{M+1}}L(x)\,\mu_{\beta}^{\gamma_{s}}(dx)$ $\leq\frac{2\hat{C}^{1/m}}{M^{1-1/m}}$ (60)
@@
$\mathbb{E}_{\hat{\pi}_{M+1}}[|X|]=\frac{\mathbb{E}_{\pi_{\gamma^{*}}}[|X|\mathds{1}_{B_{M+1}}(X)]}{\pi_{\gamma^{*}}(B_{M+1})}\leq\frac{\mathbb{E}_{\pi_{\gamma^{*}}}[|X|^{m}]^{1/m}\cdot\pi_{\gamma^{*}}(B_{M+1})^{1-1/m}}{\pi_{\gamma^{*}}(B_{M+1})}=\frac{\mathbb{E}_{\pi_{\gamma^{*}}}[|X|^{m}]^{1/m}}{\pi_{\gamma^{*}}(B_{M+1})^{1/m}}$
By using Comparison Theorem 10 and using a similar procedure in Section 2.3, under the Lyapunov condition in (26) we obtain:
$\mathbb{E}_{\pi_{\gamma^{*}}}[|X|^{m}]\leq\frac{b}{\alpha}.$ (61)
For the overflow set $B_{M+1}=\{|x|\geq(CM)^{1/m}\}$ we have by Markov’s inequality:
$\pi_{\gamma^{*}}(B_{M+1})\leq\frac{\mathbb{E}_{\pi_{\gamma^{*}}}[|X|^{m}]}{(CM)^{m/m}}\leq\frac{b/\alpha}{CM}.$
@@
$\mathbb{E}_{\hat{\pi}_{M+1}}[|X^{\prime}|]\mu_{\beta}^{\gamma_{s}}(B_{M+1})$ $\leq\frac{\mathbb{E}_{\pi_{\gamma^{*}}}[|X|^{m}]^{1/m}}{\pi_{\gamma^{*}}(B_{M+1})^{1/m}}\cdot\frac{1}{M}$
$=\Bigl(\tfrac{b}{\alpha}\Bigr)^{1/m}\Bigl(\tfrac{C}{b/\alpha}\Bigr)^{1/m}M^{-(1-1/m)}=\frac{C^{1/m}}{M^{1-1/m}}$
@@
$\int_{B_{M+1}}L(x)\,\mu_{\beta}^{\gamma_{s}}(dx)$ $\leq\frac{2C^{1/m}}{M^{1-1/m}},$
@@
$\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}[L(X)]$ $\leq\frac{2C^{1/m}}{M^{1-1/m}}\cdot\left(1-\mu_{\beta}^{\gamma_{s}}(B_{M+1})\right)+\frac{2C^{1/m}}{M^{1-1/m}}$
$\leq\frac{2C^{1/m}}{M^{1-1/m}}+\frac{2C^{1/m}}{M^{1-1/m}}=\frac{4C^{1/m}}{M^{1-1/m}}.$ (62)
@@
$\left|\hat{J}_{\beta}(x_{0})-J^{*}_{\beta}(x_{0})\right|\leq\left(\alpha_{c}+\frac{\beta\alpha_{T}\|c\|_{\infty}}{1-\beta}\right)\left(\frac{4C^{1/m}}{(M^{1-1/m})(1-\beta)}\right),$ (63)

### v2

@@ Appendix F Proof of Theorem 11
$\mu_{\beta}^{\gamma_{s}}(B_{M+1})$ $={\mu_{\beta}^{\gamma_{s}}(\{x={x_{1},\cdots,x_{n}}:\min_{i=1,\cdots,n}|x_{i}|\geq(Ck)^{1/m}\})\leq}\mathop{\rm Pr}\bigl[\|X\|_{1}\geq(Ck)^{1/m}\bigr]\leq\frac{1}{k}$ (56)
@@
$\int_{B_{M+1}}L(x)\,\mu_{\beta}^{\gamma_{s}}(dx)$ $\leq\frac{2\hat{C}^{1/m}}{k^{1-1/m}}=\frac{2\hat{C}^{1/m}}{M^{1/n(1-1/m)}}$ (57)
@@
$\mathbb{E}_{\hat{\pi}_{M+1}}[\|X\|_{1}]=\frac{\mathbb{E}_{\pi_{\gamma^{*}}}[\|X\|_{1}\mathds{1}_{B_{M+1}}(X)]}{\pi_{\gamma^{*}}(B_{M+1})}\leq\frac{\mathbb{E}_{\pi_{\gamma^{*}}}[\|X\|^{m}]^{1/m}\cdot\pi_{\gamma^{*}}(B_{M+1})^{1-1/m}}{\pi_{\gamma^{*}}(B_{M+1})}=\frac{\mathbb{E}_{\pi_{\gamma^{*}}}[\|X\|^{m}]^{1/m}}{\pi_{\gamma^{*}}(B_{M+1})^{1/m}}$
By [18, Theorem 14.2.2], under the Lyapunov condition in (25) we obtain:
$\mathbb{E}_{\pi_{\gamma^{*}}}[\|X\|_{1}^{m}]\leq\frac{b}{\alpha}.$ (58)
For the overflow set $B_{M+1}$ we have by Markov’s inequality:
$\pi_{\gamma^{*}}(B_{M+1})\leq\frac{\mathbb{E}_{\pi_{\gamma^{*}}}[\|X\|_{1}^{m}]}{(Ck)^{m/m}}\leq\frac{b/\alpha}{Ck}.$
@@
$\mathbb{E}_{\hat{\pi}_{M+1}}[\|X^{\prime}\|_{1}]\mu_{\beta}^{\gamma_{s}}(B_{M+1})\leq\frac{\mathbb{E}_{\pi_{\gamma^{*}}}[\|X\|_{1}^{m}]^{1/m}}{\pi_{\gamma^{*}}(B_{M+1})^{1/m}}\cdot\frac{1}{k}=\Bigl(\tfrac{b}{\alpha}\Bigr)^{1/m}\Bigl(\tfrac{C}{b/\alpha}\Bigr)^{1/m}k^{-(1-1/m)}=\frac{C^{1/m}}{M^{1/n(1-1/m)}}$
@@
$\int_{B_{M+1}}L(x)\,\mu_{\beta}^{\gamma_{s}}(dx)$ $\leq\frac{2C^{1/m}}{M^{1/n(1-1/m)}},$
@@
$\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}[L(X)]$ $\leq\frac{2C^{1/m}}{M^{1/n(1-1/m)}}\cdot\left(1-\mu_{\beta}^{\gamma_{s}}(B_{M+1})\right)+\frac{2C^{1/m}}{M^{1/n(1-1/m)}}$
$\leq\frac{2C^{1/m}}{M^{1/n(1-1/m)}}+\frac{2C^{1/m}}{M^{1/n(1-1/m)}}=\frac{4C^{1/m}}{M^{1/n(1-1/m)}}.$ (59)
@@
$\left|\hat{J}_{\beta}(x_{0})-J^{*}_{\beta}(x_{0})\right|\leq\left(\alpha_{c}+\frac{\beta\alpha_{T}\|c\|_{\infty}}{1-\beta}\right)\left(\frac{4C^{1/m}}{(M^{1/n(1-1/m)})(1-\beta)}\right),$ (60)